OpenStudy (anonymous):

In a certain state , fishing licenses are in the form LLL NN , where L stands for a letter of the alphabet and N stands for one digit number from 0 to 9. How many possible fishing license outcomes are there if all are allowed

OpenStudy (anonymous):

Well, there are how many letters in the alphabet?

OpenStudy (anonymous):

26 @NinjaDevo

OpenStudy (nurali):

The number of possible outcomes for "NN" are 10^2=100 The number of possible outcomes for "LLL" are 26^3=17576 The number of possible outcomes for "NNN" are 10^3=1000 Multiply them together and you get 100x17576x1000=1757600000 so there are 1757600000 possible combinations of licence plate with that particular form.

OpenStudy (anonymous):

Yup, Nurali got it. That's basically what I was going to say step-for-step. :)